A body of mass 108kg at rest is pushed a distance of 300 m by a force of 72N.calculate the time taken to cover the distance
1
Expert's answer
2019-05-27T11:04:53-0400

The body acceleration

"a=\frac{F}{m}=\frac{72\:\rm{N}}{108\:\rm{kg}}=0.67\:\rm{m/s^2}"

The distance

"d=\frac{at^2}{2}"

So, time taken to cover the distance

"t=\sqrt{\frac{2d}{a}}=\sqrt{\frac{2\times 300\:\rm{m}}{0.67\:\rm{m/s^2}}}=30\:\rm{s}"